solve each equation below by making the appropriate substitution. Be sure to check answer(s): PLEASE HELP ME!
A)X^-2+X^-1-56=0
U=
u^2=
Answer:
B)x^2/3-x^1/3-12=0
U=
U^2=
Answer:

Answer by stanbon(75887)

A)X^-2+X^-1-56=0
U=x^-1
u^2=x^-2
Answer:u^2+u-56 = 0
(u+8)(u-7)=0
u=-8 or u=7
x^-1=-8 or x^-1=7
x=-1/8 or x=1/7
--------------------------
B)x^2/3-x^1/3-12=0
U= x^(1/3)
U^2=x^(2/3)
Answer:u^2+u-12=0
(u+4)(u-3)=0
u=-4 or u=3
x^(1/3)=-4 or x^(1/3)=3
x = -64 or x=27
----------------
Cheers,
Stan H.